### v2

@@ 5.3 The rank deficient case
Previously we proved that when $f$ is locally strongly convex, super-linear convergence cannot hold for RS-RNM. Here we prove that when the Hessian $\bar{H}$ at the local optimum $\bar{x}$ is rank deficient, then RS-RNM can achieve super-linear convergence. In this whole subsection, we assume that Assumptions 5 and 5.1 are satisfied. We also denote by $r$ ($<n$) the rank of $\bar{H}$. We first prove the following lemma which is very similar to Lemma 5.14.
@@
We have, under Assumptions 5 and 5.1, that for $k$ large enough:
@@
Proof. As in the proof of Lemma 5.14, we have that
$\rho\|x_{k}-\bar{x}\|-L_{H}\|x_{k}-\bar{x}\|^{2}\leq\|\nabla f(x_{k})\|-L_{H}\|x_{k}-\bar{x}\|^{2}{\leq}\|\bar{H}(x_{k}-\bar{x})\|.$
Since $\|x_{k}-\bar{x}\|$ tends to 00, we deduce that for $k$ large enough:
@@
The other inequality is easy to deduce from (50), as in (60):
$\|g_{k}\|\leq\|\bar{H}\|\|x_{k}-\bar{x}\|+L_{H}\|x_{k}-\bar{x}\|^{2}\leq 2\lambda_{\max}(\bar{H})\|x_{k}-\bar{x}\|,$
@@
Under Assumptions 5 and 5.1. If $s\geq r$, we have that for $k$ large enough, with probability at least $1-2\exp(-s)$:
@@
Proof. By (50) applied at $k+1$, we have that
[…]
@@
$\|P_{k}g_{k+1}\|\geq\|P_{k}\bar{H}(x_{k+1}-\bar{x})\|-L_{H}\|P_{k}\|\|x_{k+1}-\bar{x}\|^{2}.$ (70)
Let $UDU^{\top}=\bar{H}$ be the diagonal decomposition of $\bar{H}$. Since $\bar{x}$ is a strict local minimizer, by Assumption 5.1, for $k$ large enough, $U$ is an orthogonal matrix independent of $P_{k}$, and hence, $\tilde{P}_{k}:=P_{k}U$ is an i.i.d. random Gaussian matrix with the same distribution as $P_{k}$. Let $y_{k+1}=U^{\top}(x_{k+1}-\bar{x})$. We have that
$\bar{H}(x_{k+1}-\bar{x})=Dy_{k+1}\quad\&\quad P_{k}\bar{H}(x_{k+1}-\bar{x})=\tilde{P}_{k}Dy_{k+1}.$ (71)
@@
where the inequality follows from Lemma 5.24. Hence, from (70) and (72), we deduce that
@@
$\|P_{k}g_{k+1}\|\geq\rho\frac{\sigma_{\min}(P_{k}^{1})}{4}\|x_{k+1}-\bar{x}\|\geq\rho\frac{\sigma_{\min}(P_{k}^{1})}{4}\frac{\|g_{k+1}\|}{2\lambda_{\max}(\bar{H})}.$
@@
Let $M\in\mathbb{R}^{n\times n}$ be any matrix. Under Assumptions 5 and 5.1, if $k$ is large enough and $s\geq r$, we have
@@
$\|P_{k}H_{k}M\|\geq\|P_{k}\bar{H}M\|-\|P_{k}(H_{k}-\bar{H})M\|.$ (73)
Let $UDU^{\top}=\bar{H}$ be the diagonal decomposition of $\bar{H}$. Similarly to the proof of Lemma 5.26, $\tilde{P}_{k}:=P_{k}U$ is an i.i.d. random Gaussian matrix with the same distribution as $P_{k}$. Let $N=U^{\top}M$. We have that $P_{k}\bar{H}M=\tilde{P}_{k}DN$. Furthermore, since $D$ has rank $r<n$, we can write $DN=\begin{pmatrix}\tilde{N}\\ 0\end{pmatrix}$, where $\tilde{N}\in\mathbb{R}^{r\times n}$. We have therefore that
$\|P_{k}\bar{H}M\|=\|{P}_{k}^{1}\tilde{N}\|,$ (74)
@@
$\|P_{k}\bar{H}M\|\geq\sigma_{\min}({P}_{k}^{1})\|\tilde{N}\|=\sigma_{\min}({P}_{k}^{1})\|DN\|=\sigma_{\min}({P}_{k}^{1})\|\bar{H}M\|,$ (75)
where the last equality holds by orthogonality of $U$. We deduce therefore, from (73) and (75) that
@@
The next lemma, similar to Lemma 5.2 of [26], is needed to control $\|\eta_{k}\|=\|c_{1}\Lambda_{k}+c_{2}\left\|g_{k}\right\|^{\gamma}\|$, where $\Lambda_{k}=\max(0,-\lambda_{\min}(P_{k}H_{k}P_{k}^{\top}))$.
@@
Under Assumption 5, for $k$ large enough, we have that with probability at least $1-2\exp(-s)$,
@@
Proof. The result is obvious when $\Lambda_{k}=0$. Let us consider the case $\Lambda_{k}>0$. Let $\lambda=(\lambda_{k}^{(1)},\dots,\lambda_{k}^{(s)})$ be a vector of eigenvalues of $P_{k}\bar{H}P_{k}^{\top}$ and we write the eigenvalue decomposition of $P_{k}\bar{H}P_{k}^{\top}$ as follows:
@@
Since $\lambda_{\min}(P_{k}H_{k}P_{k}^{\top})I_{s}-U_{k}P_{k}H_{k}P_{k}^{\top}U_{k}^{\top}$ is singular, therefore
@@
is not singular. We define
@@
$\|(\lambda_{\min}(P_{k}H_{k}P_{k}^{\top})I_{s}-diag(\lambda_{k}))^{-1}\|\leq\frac{1}{-\lambda_{\min}(P_{k}H_{k}P_{k}^{\top})}=\frac{1}{\Lambda_{k}}.$ (76)
@@
where the first inequality is a well known inequality for a singular matrix and is proved in [26, Lemma 5.1].
@@
Under Assumptions 5 and 5.1, if $s\geq r$, we have that for $k$ large enough, we have that with probability at least $1-2\exp(-s)$,
@@
Proof. Notice first that by Taylor expansion of $t\mapsto\nabla f(\bar{x}+t(x_{k}-\bar{x}))$ and by Assumption 5, we have that
$\|g_{k}-\nabla f(\bar{x})-H_{k}(x_{k}-\bar{x})\|\leq\frac{L_{H}}{2}\|x_{k}-\bar{x}\|^{2}.$ (77)
@@
=$\overset{\nabla f(\bar{x})=0}{=}\|{P_{k}^{\top}M_{k}^{-1}P_{k}(g_{k}-\nabla f(\bar{x})-H_{k}(x_{k}-\bar{x})+H_{k}(x_{k}-\bar{x}))}\|$
@@
(77)${\leq}\frac{L_{H}}{2}\|{P_{k}}\|^{2}\|{M_{k}^{-1}}\|\|{x_{k}-\bar{x}}\|^{2}+\|{P_{k}^{\top}M_{k}^{-1}P_{k}H_{k}}\|\|{x_{k}-\bar{x}}\|.$ (78)
Let us first bound the first term in the right-hand side of (5.33). When $k$ is large enough, with probability at least $1-2\exp(-s)$, we have by Lemma 2.2
$\frac{L_{H}}{2}\|{P_{k}}\|^{2}\|{M_{k}^{-1}}\|$ $\leq\frac{L_{H}}{2}\cdot\frac{\bar{\mathcal{C}}n}{s}\cdot\frac{1}{\lambda_{\min}(P_{k}H_{k}P_{k}^{\top}+c_{1}\Lambda_{k}I_{s}+c_{2}\|{g_{k}}\|^{\delta}I_{s})}$
$\leq\frac{L_{H}\bar{\mathcal{C}}n}{2c_{2}s\|{g_{k}}\|^{\delta}}$
(49)${\leq}\frac{L_{H}\bar{\mathcal{C}}n}{2c_{2}s\rho^{\delta}\|x_{k}-\bar{x}\|^{\delta}}.$
@@
$\frac{L_{H}}{2}\|{P_{k}}\|^{2}\|{M_{k}^{-1}}\|\|x_{k}-\bar{x}\|^{2}\leq\frac{L_{H}\bar{\mathcal{C}}n}{2c_{2}s\rho^{\delta}}\|x_{k}-\bar{x}\|^{2-\delta}.$ (79)
@@
$\leq 1+\frac{c_{1}\Lambda_{k}+c_{2}\|g_{k}\|^{\delta}}{(c_{1}-1)\Lambda_{k}+c_{2}\|g_{k}\|^{\delta}}$
@@
where the second inequality follows from Lemma 2.2. The results follows from (5.33) and (79) noticing that $\frac{\|x_{k}-\bar{x}\|^{2-\delta}}{\|x_{k}-\bar{x}\|}$ tends to 00, as $\delta<1$, hence for $k$ large enough
$\frac{L_{H}\bar{\mathcal{C}}n}{2c_{2}s\rho^{\delta}}\|x_{k}-\bar{x}\|^{2-\delta}\leq\frac{2}{\sigma_{\min}(P_{1}^{k})}\left(2+\frac{1}{c_{1}-1}\right)\sqrt{\frac{\bar{\mathcal{C}}n}{s}}\|x_{k}-\bar{x}\|.$
@@
Under Assumptions 5 and 5.1, for $k$ large enough and for any $s\geq r$, we have that with probability at least $1-2\exp(-s)$
$\|x_{k+1}-\bar{x}\|\leq\frac{c_{2}\Gamma}{\sigma^{2}_{\min}(P_{k}^{1})}\|x_{k}-\bar{x}\|^{1+\delta},$
@@
$\leq\frac{8\lambda_{\max}(\bar{H})}{\rho^{2}\sigma_{\min}(P^{1}_{k})}\left(\|P_{k}(g_{k+1}-g_{k}-H_{k}(x_{k+1}-x_{k}))\|+\|P_{k}g_{k}+P_{k}H_{k}(x_{k+1}-x_{k})\|\right).$
Where the first inequality holds by (49), and the second holds by Lemma 5.26. By Lemma 5.32 and ab equation similar to (77)(where $x_{k}$ is replaced by $x_{k+1}$ and $\bar{x}$ is replaced by $x_{k}$), we have that
$\|P_{k}(g_{k+1}-g_{k}-H_{k}(x_{k+1}-x_{k}))\|\leq L_{H}\|P_{k}\|\left(\frac{4}{\sigma_{\min}(P_{1}^{k})}\left(2+\frac{1}{c_{1}-1}\right)\sqrt{\frac{\bar{\mathcal{C}}n}{s}}\right)^{2}\|x_{k}-\bar{x}\|^{2}.$
Furthermore,
@@
$\frac{\eta_{k}}{\sigma_{\min}(P_{k}^{\top})}\|d_{k}\|$ $\leq\frac{1}{\sigma^{2}_{\min}(P_{k}^{\top})}\left(c_{1}\alpha\|x_{k}-\bar{x}\|^{2}+c_{2}\beta\|x_{k}-\bar{x}\|\|g_{k}\|^{\delta}\right)$
$\leq\frac{1}{\sigma^{2}_{\min}(P_{k}^{\top})}\left(c_{1}\alpha\|x_{k}-\bar{x}\|^{2}+c_{2}\beta^{\prime}\|x_{k}-\bar{x}\|^{1+\delta}\right),$
where we have used in the second inequality that $\|g_{k}\|\leq O(\|x_{k}-\bar{x}\|)$. Putting everything together, we obtain the desired result.
Notice that by using [23], we can furthermore bound $\frac{1}{\sigma_{\min}(P_{k}^{1})}$, with high probability, by $O(\frac{1}{\sqrt{s}-\sqrt{r-1}})$. Hence one can derive a version, in expectation of Theorem 5.20, where no probability bound appear, as we did for Corollary 5.12.
Finally notice that Assumption 5.1 can be easily satisfied by considering functions with low effective dimension, that is function of the form $y\mapsto f(Ry)$, where $y\in\mathbb{R}^{r}$ and where $R\in\mathbb{R}^{n\times r}$ is an unknown matrix.

### v3

@@ 5.3 The rank deficient case
Previously we proved that when $f$ is locally strongly convex, super-linear convergence cannot hold for RS-RNM. Here we prove that when the Hessian $\bar{H}$ at the local optimum $\bar{x}$ is rank deficient, then RS-RNM can achieve super-linear convergence. In this whole subsection, we assume that Section 5 and Section 5.1 are satisfied. We also denote by $r$ (<n<n) the rank of $\bar{H}$. Notice that, as a special case of $r<n$, one can consider “functions with low dimensionality” [^5] They are also called objectives with “active subspaces” [10], or “multi-ridge” [16]. [42]. For such functions, there exists a projection matrix $\Pi\in\mathbb{R}^{n\times n}$ with $\mathrm{rank}(\Pi)<n$ such that
(77) $\forall x\in\mathbb{R}^{n},~f(x)=f(\Pi x).$
Such functions are frequently encountered in many applications. For example, the loss functions of neural networks often have low rank Hessians [21, 36, 33]. This phenomenon is also prevalent in other areas such as hyper-parameter optimization for neural networks [3], heuristic algorithms for combinatorial optimization problems [23], complex engineering and physical simulation problems as in climate modeling [26], and policy search [17].
We first prove the following lemma which is very similar to Lemma 5.14.
@@
We have, under Section 5 and Section 5.1, that for $k$ large enough:
@@
Proof. As in the proof of Lemma 5.14, we have (56), i.e.,
$\rho\|x_{k}-\bar{x}\|-L_{H}\|x_{k}-\bar{x}\|^{2}{\leq}\|\bar{H}(x_{k}-\bar{x})\|.$
Since $\|x_{k}-\bar{x}\|$ tends to 0, we deduce that for $k$ large enough:
@@
The other inequality is easy to deduce from Eq. 54, as in Eq. 66:
(78) $\|g_{k}\|\leq\|\bar{H}\|\|x_{k}-\bar{x}\|+L_{H}\|x_{k}-\bar{x}\|^{2}\leq 2\lambda_{\max}(\bar{H})\|x_{k}-\bar{x}\|,$
@@
Under Section 5 and Section 5.1. If $s\geq r$, we have that for $k$ large enough, with probability at least $1-2\exp(-s)$:
@@
Proof. By Eq. 54 applied at $k+1$, we have that
$\nabla f(x_{k+1})=\int_{0}^{1}\nabla^{2}f(\bar{x}+t(x_{k+1}-\bar{x}))(x_{k+1}-\bar{x})dt.$
@@
(79) $\|P_{k}g_{k+1}\|\geq\|P_{k}\bar{H}(x_{k+1}-\bar{x})\|-L_{H}\|P_{k}\|\|x_{k+1}-\bar{x}\|^{2}.$
Let $UDU^{\top}=\bar{H}$ be the diagonal decomposition of $\bar{H}$. Since $\bar{x}$ is a strict local minimizer, by Section 5.1, for $k$ large enough, $U$ is an orthogonal matrix independent of $P_{k}$, and hence, $\tilde{P}_{k}:=P_{k}U$ is an i.i.d. random Gaussian matrix with the same distribution as $P_{k}$. Let $y_{k+1}=U^{\top}(x_{k+1}-\bar{x})$. We have that
(80) $\bar{H}(x_{k+1}-\bar{x})=UDy_{k+1}\quad\mbox{ and thus, }\quad P_{k}\bar{H}(x_{k+1}-\bar{x})=\tilde{P}_{k}Dy_{k+1}.$
@@
where the inequality follows from Lemma 5.24. Hence, from Eq. 79 and Eq. 81, we deduce that
@@
[…]
@@
Let $M\in\mathbb{R}^{n\times n}$ be any matrix. Under Section 5 and Section 5.1, if $k$ is large enough and $s\geq r$, we have
@@
(82) $\|P_{k}H_{k}M\|\geq\|P_{k}\bar{H}M\|-\|P_{k}(H_{k}-\bar{H})M\|.$
Let $UDU^{\top}=\bar{H}$ be the diagonal decomposition of $\bar{H}$. Similarly to the proof of Lemma 5.26, $\tilde{P}_{k}:=P_{k}U$ is an i.i.d. random Gaussian matrix with the same distribution as $P_{k}$. Using $N:=U^{\top}M$, we have that $P_{k}\bar{H}M=\tilde{P}_{k}DN$. Furthermore, since $D$ has rank $r<n$, we can write $DN=\begin{pmatrix}\tilde{N}\\ 0\end{pmatrix}$, where $\tilde{N}\in\mathbb{R}^{r\times n}$. We have therefore that
(83) $\|P_{k}\bar{H}M\|=\|{P}_{k}^{1}\tilde{N}\|$,
@@
(84) $\|P_{k}\bar{H}M\|\geq\sigma_{\min}({P}_{k}^{1})\|\tilde{N}\|=\sigma_{\min}({P}_{k}^{1})\|DN\|=\sigma_{\min}({P}_{k}^{1})\|\bar{H}M\|,$
where the last equality holds by orthogonality of $U$. We deduce therefore, from Eq. 82 and Eq. 84 that
@@
The next lemma, similar to Lemma 5.2 of [39], is needed to control $\eta_{k}=c_{1}\Lambda_{k}+c_{2}\left\|g_{k}\right\|^{\gamma}$, where $\Lambda_{k}=\max(0,-\lambda_{\min}(P_{k}H_{k}P_{k}^{\top}))$.
@@
Under Section 5, for $k$ large enough, we have that with probability at least $1-2\exp(-s)$,
@@
Proof. The result is obvious when $\Lambda_{k}=0$. Let us consider the case $\Lambda_{k}>0$. Let $\lambda_{k}=(\lambda_{k}^{(1)},\dots,\lambda_{k}^{(s)})$ be a vector of eigenvalues of $P_{k}\bar{H}P_{k}^{\top}$ and we write the eigenvalue decomposition of $P_{k}\bar{H}P_{k}^{\top}$ as follows:
@@
Notice that $\lambda_{\min}(P_{k}H_{k}P_{k}^{\top})I_{s}-U_{k}P_{k}H_{k}P_{k}^{\top}U_{k}^{\top}$ is singular. Furthermore,
@@
is not singular as $\lambda_{\min}(P_{k}H_{k}P_{k}^{\top})<0$ by assumption and $diag(\lambda_{k})$ is positive. We define
@@
(85) $\|(\lambda_{\min}(P_{k}H_{k}P_{k}^{\top})I_{s}-diag(\lambda_{k}))^{-1}\|\leq\frac{1}{-\lambda_{\min}(P_{k}H_{k}P_{k}^{\top})}=\frac{1}{\Lambda_{k}}.$
@@
[…]
where the first inequality is a well known inequality for a singular matrix and is proved in [39, Lemma 5.1].
@@
Under Section 5 and Section 5.1, if $s\geq r$, we have that for $k$ large enough, we have that with probability at least $1-2\exp(-s)$,
@@
Proof. Notice first that by Taylor expansion of $t\mapsto\nabla f(\bar{x}+t(x_{k}-\bar{x}))$ and by Section 5, we have that
(86) $\|g_{k}-\nabla f(\bar{x})-H_{k}(x_{k}-\bar{x})\|\leq\frac{L_{H}}{2}\|x_{k}-\bar{x}\|^{2}.$
@@
$\overset{\nabla f(\bar{x})=0}{=}\|{P_{k}^{\top}M_{k}^{-1}P_{k}(g_{k}-\nabla f(\bar{x})-H_{k}(x_{k}-\bar{x})+H_{k}(x_{k}-\bar{x}))}\|$
@@
[…]
Let us first bound the first term in the right-hand side of 5.33. When $k$ is large enough, with probability at least $1-2\exp(-s)$, we have by Lemma 2.2
$\frac{L_{H}}{2}\|{P_{k}}\|^{2}\|{M_{k}^{-1}}\|$ $\leq\frac{L_{H}}{2}\cdot\frac{\bar{\mathcal{C}}n}{s}\cdot\frac{1}{\lambda_{\min}(P_{k}H_{k}P_{k}^{\top}+c_{1}\Lambda_{k}I_{s}+c_{2}\|{g_{k}}\|^{\gamma}I_{s})}$
$\leq\frac{L_{H}\bar{\mathcal{C}}n}{2c_{2}s\|{g_{k}}\|^{\gamma}}$
[…]
@@
(88) $\frac{L_{H}}{2}\|{P_{k}}\|^{2}\|{M_{k}^{-1}}\|\|x_{k}-\bar{x}\|^{2}\leq\frac{L_{H}\bar{\mathcal{C}}n}{2c_{2}s\rho^{\gamma}}\|x_{k}-\bar{x}\|^{2-\gamma}.$
@@
$\leq 1+\frac{c_{1}\Lambda_{k}+c_{2}\|g_{k}\|^{\gamma}}{(c_{1}-1)\Lambda_{k}+c_{2}\|g_{k}\|^{\gamma}}$
@@
where the second inequality follows from Lemma 2.2. The results follows from 5.33 and Eq. 88 noticing that $\frac{\|x_{k}-\bar{x}\|^{2-\gamma}}{\|x_{k}-\bar{x}\|}$ tends to 0, as $\gamma<1$, hence for $k$ large enough
$\frac{L_{H}\bar{\mathcal{C}}n}{2c_{2}s\rho^{\gamma}}\|x_{k}-\bar{x}\|^{2-\gamma}\leq\frac{2}{\sigma_{\min}(P_{1}^{k})}\left(2+\frac{1}{c_{1}-1}\right)\sqrt{\frac{\bar{\mathcal{C}}n}{s}}\|x_{k}-\bar{x}\|.$
@@
Under Section 5 and Section 5.1, for $k$ large enough and for any $s\geq r$, we have that with probability at least $1-2\exp(-s)$
$\|x_{k+1}-\bar{x}\|\leq\frac{c_{2}\Gamma}{\sigma^{2}_{\min}(P_{k}^{1})}\|x_{k}-\bar{x}\|^{1+\gamma},$
@@
(89) $\leq\frac{8\lambda_{\max}(\bar{H})}{\rho^{2}\sigma_{\min}(P^{1}_{k})}\left(\|P_{k}(g_{k+1}-g_{k}-H_{k}(x_{k+1}-x_{k}))\|+\|P_{k}g_{k}+P_{k}H_{k}(x_{k+1}-x_{k})\|\right),$
where the first inequality holds by Eq. 53, and the second holds by Lemma 5.26.
By Lemma 5.32 and an equation similar to Eq. 86 (where $x_{k}$ is replaced by $x_{k+1}$ and $\bar{x}$ is replaced by $x_{k}$), we have that
(90) $\|P_{k}(g_{k+1}-g_{k}-H_{k}(x_{k+1}-x_{k}))\|\leq L_{H}\|P_{k}\|\left(\frac{4}{\sigma_{\min}(P_{1}^{k})}\left(2+\frac{1}{c_{1}-1}\right)\sqrt{\frac{\bar{\mathcal{C}}n}{s}}\right)^{2}\|x_{k}-\bar{x}\|^{2}.$
From the updated rule $x_{k+1}=x_{k}-t_{k}P_{k}^{\mathsf{T}}M_{k}^{-1}P_{k}g_{k}$ in Algorithm 1, we see that $x_{k+1}-x_{k}=-t_{k}P_{k}^{\mathsf{T}}M_{k}^{-1}P_{k}g_{k}$. From now on, we will show that $t_{k}=1$ for $k$ large enough. Indeed by (4.12), we have that
$f(x_{k})-f(x_{k}+t_{k}^{\prime}d_{k})+\alpha t_{k}^{\prime}g_{k}^{\mathsf{T}}d_{k}\geq\frac{\bar{\mathcal{C}}n}{2s}L_{H}{t_{k}^{\prime}}^{2}\left\|d_{k}\right\|\left(\frac{c_{2}s\left\|g_{k}\right\|^{\gamma}}{\bar{\mathcal{C}}L_{H}n\left\|d_{k}\right\|}-t_{k}^{\prime}\right)\left\|M_{k}^{-1}P_{k}g_{k}\right\|^{2}.$
Hence, by Section 5.1 and Lemma 5.32, we deduce that there exists some constant $\mathcal{C}_{1}$ such that
$f(x_{k})-f(x_{k}+t_{k}^{\prime}d_{k})+\alpha t_{k}^{\prime}g_{k}^{\mathsf{T}}d_{k}\geq\frac{\bar{\mathcal{C}}n}{2s}L_{H}{t_{k}^{\prime}}^{2}\left\|d_{k}\right\|\left(\frac{\mathcal{C}_{1}}{\|x_{k}-\bar{x}\|^{1-\gamma}}-t_{k}^{\prime}\right)\left\|M_{k}^{-1}P_{k}g_{k}\right\|^{2},$
proving that we can take $t^{\prime}_{k}=1$ if $\|x_{k}-\bar{x}\|$ is small enough.
Now notice that for $k$ large enough, $t_{k}=1$, hence
@@
$\frac{\eta_{k}}{\sigma_{\min}(P_{k}^{\top})}\|d_{k}\|$ $\leq\frac{1}{\sigma^{2}_{\min}(P_{k}^{\top})}\left(c_{1}\alpha\|x_{k}-\bar{x}\|^{2}+c_{2}\beta\|x_{k}-\bar{x}\|\|g_{k}\|^{\gamma}\right)$
$\leq\frac{1}{\sigma^{2}_{\min}(P_{k}^{\top})}\left(c_{1}\alpha\|x_{k}-\bar{x}\|^{2}+c_{2}\beta^{\prime}\|x_{k}-\bar{x}\|^{1+\gamma}\right),$
where we have used in the second inequality that $\|g_{k}\|\leq O(\|x_{k}-\bar{x}\|)$. Now by (5.35), (90) and the above, we obtain the desired result.
Notice that by using [35], we can furthermore bound $\frac{1}{\sigma_{\min}(P_{k}^{1})}$, with high probability, by $O(\frac{1}{\sqrt{s}-\sqrt{r-1}})$.
Let us consider a function with low dimensionality, i.e. satisfying (77). Let us write $\Pi=R^{\top}R$, where $R\in\mathbb{R}^{s\times n}$ and let us define $g:y\in\mathbb{R}^{s}\mapsto f(R^{\top})$. Hence, we have that $g(Rx)=f(\Pi x)=f(x)$. By denoting $y_{k}:=Rx_{k}\in\mathbb{R}^{s}$ and assuming that the function $g(y)$ is strongly convex, locally near $\bar{y}:=R\bar{x}$, it is easy to see that Section 5.1 is satisfied for the sequence $\{y_{k}\}$, locally, i.e., there exists $\rho>0$ such that for $k$ large enough;
$\|\nabla g(y_{k})\|\geq\rho\|y_{k}-\bar{y}\|$
holds. Hence, we can prove that there exists some constant $\mathcal{K}>$ such that the following inequality holds with high probability.
$\|y_{k+1}-\bar{y}\|\leq\mathcal{K}\|y_{k}-\bar{y}\|^{1+\gamma}.$
By strong convexity of $g(y)$, we know that there exists two constant $l_{1}>l_{2}>0$ such that
$l_{2}(g(y_{k+1}-g(\bar{y})))\leq\|y_{k+1}-\bar{y}\|\leq l_{1}(g(y_{k+1}-g(\bar{y}))).$
Hence by following the same proof as in Corollary 5.12, we can obtain the following super-linear rate in expectation:
Theorem 5.36.
Assume that there exists a function $g:y\in\mathbb{R}^{s}\mapsto g(y)$ such that $g(Rx)=f(x)$, for some matrix $R\in\mathbb{R}^{s\times n}$ ($s<n$). If the function $g(y)$ is strongly convex, locally near $R\bar{x}$, then there exists a constant $\mathcal{K}^{\prime}>0$, such that if $k$ is large enough:
(91) $\mathbb{E}\left[f(x_{k+1})-f(\bar{x})\right]\leq\mathcal{K}^{\prime}\mathbb{E}\left[f(x_{k})-f(\bar{x})\right]^{1+\gamma},$